OpenStudy (anonymous):

According to this property, 6a + a = 6a + 1a. what property is it

OpenStudy (anonymous):

Distributive property for me :)

OpenStudy (swissgirl):

Nah I dont think so

OpenStudy (swissgirl):

Its Identity or rather known as Multiplicative Identity

OpenStudy (swissgirl):

1*a=a*1=a

OpenStudy (swissgirl):

And additive identity is a+0=0+a=a

OpenStudy (anonymous):

It's commutative property xD

OpenStudy (swissgirl):

Commutative is 6a+a=a+6a So nope it aint commutative

OpenStudy (anonymous):

addition is commutative xD